# MIB: A Mechanistic Interpretability Benchmark

### v1

@@ Abstract
How can we know whether new mechanistic interpretability methods achieve real improvements? In pursuit of meaningful and lasting evaluation standards, we propose MIB, a Mechanistic Interpretability Benchmark with two tracks spanning four tasks and five models. MIB favors methods that precisely and concisely recover relevant causal pathways or specific causal variables in neural language models. The circuit localization track compares methods that locate the model components—and connections between them—most important for performing a task (e.g., attribution patching or information flow routes). The causal variable localization track compares methods that featurize a hidden vector, e.g., sparse autoencoders (SAE) or distributed alignment search (DAS), and locate model features for a causal variable relevant to the task. Using MIB, we find that attribution and mask optimization methods perform best on circuit localization. For causal variable localization, we find that the supervised DAS method performs best, while SAEs features are not better than neurons, i.e., standard dimensions of hidden vectors. These findings illustrate that MIB enables meaningful comparisons of methods, and increases our confidence that there has been real progress in the field.
@@ 2.1 Tasks
The number of instances in each dataset and split is summarized in Table 4 (App. B). Each task comes with a training split on which users can discover circuits or causal variables, and a validation split on which users can tune their methods or hyperparameters. We also create two test sets per task: public and private. The public test set enables faster iteration on methods. We release the train, validation, and public test sets on Huggingface. The private test set is not visible to users; they must upload either their circuits or their locations and featurizers to an API (see §2.4).
@@
The indirect object identification (IOI) task, first proposed by Wang et al. (2023), is one of the most studied tasks in MI. IOI has sentences like “When Mary and John went to the store, John gave an apple to _”, containing a subject (“John”) and an indirect object (“Mary”), which should be completed with the indirect object. Even small LMs can achieve high accuracy; thus, it has been well studied (Huben et al., 2024; Conmy et al., 2023; Merullo et al., 2024). We generate 40,000 instances. We ensure that each name tokenizes to a single token for all models we test. To assess generalization, the private test set contains names and direct objects that are not contained in the public train or test set. See App. B.1 for details.
@@
Math-related tasks are common in MI (Stolfo et al., 2023; Nanda et al., 2023; Zhang et al., 2024; Nikankin et al., 2025) and interpretability research more broadly (Liu et al., 2023; Huang et al., 2024b). We follow Stolfo et al. in defining the task as performing operations with two operands of up to two digits each. Given a pair of numbers and an operator, the model must predict the result of the operation, e.g., “What is the sum of 13 and 25?”. To create the dataset, we enumerate all possible pairs of one-digit and two-digit numbers and generate queries for addition and subtraction, yielding about 75,000 instances. Following Karpas et al. (2022) and Stolfo et al. (2023), we use six natural language templates for each operand pair to ensure we are isolating robust behavior. See App. B.2 for details.
@@
Finally, we analyze the ARC dataset (Clark et al., 2018), which comprises grade-school-level multiple-choice science questions. This is a representative task for evaluating basic scientific knowledge in LMs (Brown et al., 2020; Jiang et al., 2023; Dubey et al., 2024). Our work presents the first causal investigation of LM performance on this dataset, and among the first MI investigations of a more realistic task used to benchmark state-of-the-art LMs. We follow the dataset’s original partition to Easy and Challenge subsets (5000 and 2500 instances, respectively) and analyze them separately; this is due to a large accuracy difference on the two subsets (Table 1). We maintain the original 4-choice multiple-choice prompt formatting (see App. B.4), making this dataset related in format to, but more challenging than, MCQA.
@@ 2.2 Counterfactual Inputs
We provide counterfactual inputs for each instance in the train, validation, and test sets, where the mappings from the original inputs to the counterfactual inputs are fixed to ensure consistency in evaluation. These are provided on Huggingface. See App. B, Tables 5, 7, 8, and 9 for examples and App. B for more details.
@@ 2.3 Models
Table 1: Model performance (0-shot, greedy generation) for all models and tasks on our public test splits. For results using ranked-choice scoring and more details on evaluation, see App. B.5.
@@
Llama-3.1 8B 0.69 0.92 0.95 0.89 0.93 0.79
Gemma-2 2B 0.82 1.00 0.74 0.62 0.79 0.59
Qwen-2.5 0.5B 0.99 1.00 0.30 0.33 0.73 0.58
GPT2-Small 0.92 0.06 0.00 0.09 0.03 0.03
@@
For a mechanistic analysis to be meaningful for a given model and task, the model generally should be able to perform the task well (and perform well on counterfactuals). Performance for all models and tasks using 0-shot prompts [^5] In-context learning is itself a behavior worth studying mechanistically, but one that is challenging to disentangle from task-specific behavior, leading many MI studies to use 0-shot prompts. and greedy decoding is in Table 1; performance on counterfactual inputs are in App. B.5. Gemma 2 and Llama are generally capable across the board, performing well on MCQA and ARC (Easy). GPT-2 Small has been extensively studied in MI; it is significantly smaller than the other models we investigate and less capable (and thus may rely on qualitatively different mechanisms than larger models that have been exposed to more data). Qwen-2.5 performs well relative to its size but has not yet been extensively studied.
@@ 2.4 Leaderboard
We have constructed two online leaderboards (one for each track) hosted on Huggingface to receive user submissions and display results. We intend for the leaderboards to serve as a living public artifact that both incentivizes progress in MI and advises users of MI tools on the state of the art. More details about the leaderboards, including screenshots, are in App. B.6.
@@ 3.1 Circuit Metrics
In a realistic neural network, it is difficult to anticipate the best-case and worst-case CPR or CMD values, meaning we cannot bound the metric without losing information. If we knew the ground-truth circuit, we could instead compute precision and recall. Thus, inspired by InterpBench (Gupta et al., 2024), we train a model that implements a known ground-truth circuit for IOI. Because we know which edges are in the circuit, we report AUROC ([0,1]($\in$[0,1]( ∈ [ 0 , 1 ]) over edges. See App. D for InterpBench model training and implementation details.
@@
Table 2: CMD scores across circuit localization methods and ablation types (lower is better), and AUROC scores for InterpBench (higher is better). All evaluations were performed using counterfactual ablations. Arithmetic scores are averaged across addition and subtraction; see Table 16 (App. C.3) for separate scores. We bold and underline the best and second-best methods per column, respectively.
@@ 3.2 Circuit Localization Baselines
Here, we evaluate common circuit localization methods. We evaluate each model [^9] We do not evaluate Gemma or Qwen on Arithmetic, as they tokenize numbers such that each digit has its own token. and method [^10] Exact activation patching and optimal ablations become intractable with respect to runtime as models scale. UGS becomes intractable with respect to memory requirements as models scale. in §2 where possible. We compare across multiple axes of variation, including (a) circuit localization method (see below), (b) granularity, including edge-level and neuron-level circuits, and (c) ablation type, including counterfactual (CF) ablations, mean ablations, and “optimal ablations” (OA; Li & Janson, 2024). For all methods, we assign importance scores to all edges or nodes, and either include the top-scoring components or perform greedy search; see App. C.1 for details.
@@
Unfortunately, AP approximates IE poorly (Syed et al., 2024). AP with Integrated Gradients (Sundararajan et al., 2017) or AP-IG, improves on AP by performing multiple steps of AP, trading off speed for approximation quality. We test two AP-IG variants: (i) AP-IG-inputs (Hanna et al., 2024) and (ii) AP-IG-activations (Marks et al., 2025). AP-IG-inputs requires $O(Z)$ forward passes, while AP-IG-activations requires $O(Z\cdot L)$, where $Z$ is the number of AP steps, and $L$ is the number of model layers. We use $Z=5$, following Hanna et al. (2024). We explore AP-IG at the node level (NAP-IG) and edge level (EAP-IG). See App. C.2 for definitions and details.
Information flow routes (IFR; Ferrando & Voita, 2024) is a non-counterfactual-based and non-causal method that includes edge $(u,v)$ in $\mathcal{C}$ if the output vector of $u$ and input vector of $v$ are highly similar; this is taken as evidence of a writing operation. See App. C.2 for details on how we adapted IFR to our formalization of circuits.
@@
As a mask-based method, we employ Uniform Gradient Sampling (UGS; Li & Janson, 2024), which uses CF ablations. Li & Janson also propose optimal ablations (OA), in which the mask is learned jointly with the value used to ablate non-circuit edges; they prefer OA as it is both independent of the example being ablated (unlike CF ablations) and minimally harmful (unlike mean ablations). Due to computational constraints, we do not do this, but instead learn OA vectors by optimizing them to minimize the expected cross-entropy loss on the task dataset. We then use these vectors as ablation values when running circuit localization with EAP. See App. C.2 for details.
@@ 3.3 Results
We present CMD and AUROC scores (Table 2) for each method and task where it is tractable to run; see Table 13 in App. C for CPR scores. On both metrics, EAP-IG-inputs with CF ablations generally performs best. EAP-IG-activations and UGS are competitive with EAP-IG-inputs w.r.t. CMD, but have higher runtime and memory usage respectively. For CPR, they are less competitive; this is unsurprising for UGS, as it directly optimizes for maintaining model behavior, rather than for finding solely positively impactful components.
@@ 4.2 Causal Variable Localization Baselines
We consider five baselines for constructing the featurizer $\mathcal{F}$ and selecting the features $\Pi_{X}$. See App. E for further details.
@@ Results.
We target the residual stream of the last token of the input and the correct choice letter token at each layer. We generally see strong evidence (Figures 3(c) and 3(e); App. E.5.1) of the causal model $\mathcal{H}_{\mathbf{MCQA}}$ being a faithful abstraction, with DAS successfully disentangling the ordering ID variable $X_{\text{Order}}$ from the output token $O_{\text{Answer}}$ in many layers. Even the full vector baseline successfully localizes the variable in some layers, though it performs poorly on average because both variables are aligned with the same features.
Table 3: Baseline results for the causal variable localization track. In each table, the first row is the task, the second row is the model, and the third row is the causal variable. For Arithmetic, MCQA, and ARC (Easy), we report interchange intervention accuracy, i.e., the proportion of aligned interventions on the causal model and deep learning model that result in the same output token(s); higher is better. For each method of aligning a causal variable to LM features, we report the mean across counterfactual datasets and layers in the low-level model. In parenthesis and bold, we report the best alignment across all layers. For IOI, we report the mean-squared error between the causal model logit and the deep learning model logit; lower is better. See App. E.5 for more detailed results by task.
@@
For each baseline, we target the last token and the last token of the second operand. We see poor performance across the board (Figure 3(d); App. E.5.2). The Gemma-2 results are at chance and the Llama-3.1 results have only faint signs of success. The difference in results is likely due to only Llama-3.1 tokenizing multi-digit numbers as single tokens. There may be no “carry-the-one” variable present in these models. Alternatively, the variable might be represented in a non-linear features space, e.g., an onion representation (Csordás et al., 2024). Future submissions that beat baselines will require genuine progress.
@@
For each baseline, we target the last token and the last token of the city entity in the prompt. We generally see evidence that the attributes can be localized and disentangled (Figure 3(g); App. E.5.3), though the Llama-3 results are weaker. Because the dataset requires balancing causing an attribute to change with not changing the other attributes, the full vector baseline entirely fails completely.
@@
First, we replicate these experiments on our curated datasets and fit our own linear model that we use to define a causal model $\mathcal{H}_{\mathbf{IOI}}$ that predicts the logit difference between the indirect object and the subject. The high-level causal model takes in a text input $T$ and computes the token and positional information $S_{\text{Tok}}$ and $S_{\text{Pos}}$ encoding the subject token identity and position, respectively. Then, the output variable $O$: (i) checks if the token and position variables $S_{\text{Tok}}$ and $S_{\text{Pos}}$ match the input $T$ and inverts the signal if a mismatch is detected, and (ii) computes the logit difference between the indirect object and subject as a linear function of these binary variables.
@@
Broadly, we find evidence that the position and token variable can be disentangled (Figure 3(f); App. E.5.4). For the full vector baseline, we conduct a brute force search and find an alignment of position to heads 7.3, 7.9 ($\Pi_{S_{\text{Pos}}}$), and 8.6 and token to head 8.10 ($\Pi_{S_{\text{Tok}}}$), to be better than other alignments to entire heads. While the variables can be disentangled at the level of heads, even better results are achieved when each variable is aligned to features of heads.
@@ 4.7 General Discussion
DAS is the only method that learns features with supervision from the high-level causal model, so it is not surprising this method performs best.
DBM on standard hidden dimensions is successful, while DBM with PCA is better in some cases.
The success of DBM shows that the dimensions of untransformed hidden vectors can be useful units of analysis; however, DAS outperforming DBM shows that non-basis-aligned directions in activation space are generally better units of analysis. Principal components also provide an improved unit of analysis when aligning the ordering variable $X_{\text{Order}}$ to the Llama-3.1 model for the MCQA task or the position $S_{\text{Pos}}$ and token $S_{\text{Tok}}$ variables in GPT-2 small.
DBM on SAE features consistently fail to provide a better unit of analysis than standard dimensions.
Across the board, SAE features are the worst performing method other than the full vector baselines with two exceptions: (1) the answer variable $O_{\text{Answer}}$ in the MCQA task where PCA performs worse and (2) the continent variable in the Gemma-2 model where PCA and DBM on standard dimensions perform one point worse. We suspect the poor performance of PCA on MCQA is due to the low-data regime of $\approx$100 examples harming the quality of the principal components we compute. This is in line with the results from the steering benchmark AxBench (Wu et al., 2025), in which SAEs struggle against simple steering baselines.
@@
There are several clear opportunities for future submissions to the causal variable track to improve upon our baselines. First, our baselines provide weak evidence at best for the carry-the-one variable of the two digit addition task. Second, while we only run baselines for IOI on the GPT-2 attention heads identified by (Wang et al., 2023), we encourage future submissions to locate and featurize attention heads in the other three models. Third, we did not conduct exhaustive hyperparameter searches for any of the baseline methods, so all tasks likely have room for improving the best alignment and average alignment across layers.

### v2

@@ Abstract
How can we know whether new mechanistic interpretability methods achieve real improvements? In pursuit of lasting evaluation standards, we propose MIB, a Mechanistic Interpretability Benchmark, with two tracks spanning four tasks and five models. MIB favors methods that precisely and concisely recover relevant causal pathways or causal variables in neural language models. The circuit localization track compares methods that locate the model components—and connections between them—most important for performing a task (e.g., attribution patching or information flow routes). The causal variable localization track compares methods that featurize a hidden vector, e.g., sparse autoencoders (SAEs) or distributed alignment search (DAS), and align those features to a task-relevant causal variable. Using MIB, we find that attribution and mask optimization methods perform best on circuit localization. For causal variable localization, we find that the supervised DAS method performs best, while SAE features are not better than neurons, i.e., non-featurized hidden vectors. These findings illustrate that MIB enables meaningful comparisons, and increases our confidence that there has been real progress in the field.
@@ 2.1 Tasks
The number of instances in each dataset and split is summarized in Table 5 (App. C). Each task comes with a training split on which users can discover circuits or causal variables, and a validation split on which users can tune their methods or hyperparameters. We also create two test sets per task: public and private. The public test set enables faster iteration on methods. We release the train, validation, and public test sets on Huggingface. The private test set is not visible to users; they must upload either their circuits or their locations and featurizers to an API (see §2.4).
@@
The indirect object identification (IOI) task, first proposed by Wang et al. (2023), is one of the most studied tasks in MI. IOI has sentences like “When Mary and John went to the store, John gave an apple to _”, containing a subject (“John”) and an indirect object (“Mary”), which should be completed with the indirect object. Even small LMs can achieve high accuracy; thus, it has been well studied (Huben et al., 2024; Conmy et al., 2023; Merullo et al., 2024). We generate 40,000 instances. We ensure that each name tokenizes to a single token for all models we test. To assess generalization, the private test set contains names and direct objects that are not contained in the public train or test set. See App. C.1 for details.
@@
Math-related tasks are common in MI (Stolfo et al., 2023; Nanda et al., 2023; Zhang et al., 2024; Nikankin et al., 2025) and interpretability research more broadly (Liu et al., 2023; Huang et al., 2024b). We follow Stolfo et al. in defining the task as performing operations with two operands of up to two digits each. Given a pair of numbers and an operator, the model must predict the result of the operation, e.g., “What is the sum of 13 and 25?”. To create the dataset, we enumerate all possible pairs of one-digit and two-digit numbers and generate queries for addition and subtraction, yielding about 75,000 instances. Following Karpas et al. (2022) and Stolfo et al. (2023), we use six natural language templates for each operand pair to ensure we are isolating robust behavior. See App. C.2 for details.
@@
Finally, we analyze the ARC dataset (Clark et al., 2018), which comprises grade-school-level multiple-choice science questions. This is a representative task for evaluating basic scientific knowledge in LMs (Brown et al., 2020; Jiang et al., 2023; Dubey et al., 2024). Our work presents the first causal investigation of LM performance on this dataset, and among the first MI investigations of a more realistic task used to benchmark state-of-the-art LMs. We follow the dataset’s original partition to Easy and Challenge subsets (5000 and 2500 instances, respectively) and analyze them separately; this is due to a large accuracy difference on the two subsets (Table 1). We maintain the original 4-choice multiple-choice prompt formatting (see App. C.4), making this dataset related in format to, but more challenging than, MCQA.
@@ 2.2 Counterfactual Inputs
We provide counterfactual inputs for each instance in the train, validation, and test sets, where the mappings from the original inputs to the counterfactual inputs are fixed to ensure consistency in evaluation. These are provided on Huggingface. See App. C, Tables 6, 8, 9, and 10 for examples and App. C for more details.
@@ 2.3 Models
Table 1: Model performance (0-shot, greedy generation) for all models and tasks on our public test splits. For results using ranked-choice scoring and more details on evaluation, see App. C.5.
@@
Llama-3.1 8B 0.71 0.92 0.96 0.88 0.93 0.79
Gemma-2 2B 0.83 1.00 0.65 0.43 0.79 0.59
Qwen-2.5 0.5B 0.99 1.00 0.37 0.29 0.73 0.58
GPT2-Small 0.92 0.06 0.00 0.10 0.03 0.03
@@
For a mechanistic analysis to be meaningful for a given model and task, the model generally should be able to perform the task well (and perform well on counterfactuals). Performance for all models and tasks using 0-shot prompts [^5] In-context learning is itself a behavior worth studying mechanistically, but one that is challenging to disentangle from task-specific behavior, leading many MI studies to use 0-shot prompts. and greedy decoding is in Table 1; performance on counterfactual inputs are in App. C.5. Gemma 2 and Llama are generally capable across the board, performing well on MCQA and ARC (Easy). GPT-2 Small has been extensively studied in MI; it is significantly smaller than the other models we investigate and less capable (and thus may rely on qualitatively different mechanisms than larger models that have been exposed to more data). Qwen-2.5 performs well relative to its size but has not yet been extensively studied.
@@ 2.4 Leaderboard
We have constructed two online leaderboards (one for each track) hosted on Huggingface to receive user submissions and display results. We intend for the leaderboards to serve as a living public artifact that both incentivizes progress in MI and advises users of MI tools on the state of the art. More details about the leaderboards, including screenshots, are in App. C.6.
@@ 3.1 Circuit Metrics
In a realistic neural network, it is difficult to anticipate the best-case and worst-case CPR or CMD values, meaning we cannot bound the metric without losing information. If we knew the ground-truth circuit, we could instead compute precision and recall. Thus, inspired by InterpBench (Gupta et al., 2024), we train a model that implements a known ground-truth circuit for IOI. Because we know which edges are in the circuit, we report AUROC ([0,1]($\in$[0,1]( ∈ [ 0 , 1 ]) over edges. See App. E for InterpBench model training and implementation details.
@@
Table 2: CMD scores across circuit localization methods and ablation types (lower is better), and AUROC scores for InterpBench (higher is better). All evaluations were performed using counterfactual ablations. Arithmetic scores are averaged across addition and subtraction; see Table 17 (App. D.3) for separate scores. We bold and underline the best and second-best methods per column, respectively.
@@ 3.2 Circuit Localization Baselines
Here, we evaluate common circuit localization methods. We evaluate each model [^9] We do not evaluate Gemma or Qwen on Arithmetic, as they tokenize numbers such that each digit has its own token. and method [^10] Exact activation patching and optimal ablations become intractable with respect to runtime as models scale. UGS becomes intractable with respect to memory requirements as models scale. in §2 where possible. We compare across multiple axes of variation, including (a) circuit localization method (see below), (b) granularity, including edge-level and neuron-level circuits, and (c) ablation type, including counterfactual (CF) ablations, mean ablations, and “optimal ablations” (OA; Li & Janson, 2024). For all methods, we assign importance scores to all edges or nodes, and either include the top-scoring components or perform greedy search; see App. D.1 for details.
@@
Unfortunately, AP approximates IE poorly (Syed et al., 2024). AP with Integrated Gradients (Sundararajan et al., 2017) or AP-IG, improves on AP by performing multiple steps of AP, trading off speed for approximation quality. We test two AP-IG variants: (i) AP-IG-inputs (Hanna et al., 2024) and (ii) AP-IG-activations (Marks et al., 2025). AP-IG-inputs requires $O(Z)$ forward passes, while AP-IG-activations requires $O(Z\cdot L)$, where $Z$ is the number of AP steps, and $L$ is the number of model layers. We use $Z=5$, following Hanna et al. (2024). We explore AP-IG at the node level (NAP-IG) and edge level (EAP-IG). See App. D.2 for definitions and details.
Information flow routes (IFR; Ferrando & Voita, 2024) is a non-counterfactual-based and non-causal method that includes edge $(u,v)$ in $\mathcal{C}$ if the output vector of $u$ and input vector of $v$ are highly similar; this is taken as evidence of a writing operation. See App. D.2 for details on how we adapted IFR to our formalization of circuits.
@@
As a mask-based method, we employ Uniform Gradient Sampling (UGS; Li & Janson, 2024), which uses CF ablations. Li & Janson also propose optimal ablations (OA), in which the mask is learned jointly with the value used to ablate non-circuit edges; they prefer OA as it is both independent of the example being ablated (unlike CF ablations) and minimally harmful (unlike mean ablations). Due to computational constraints, we do not do this, but instead learn OA vectors by optimizing them to minimize the expected cross-entropy loss on the task dataset. We then use these vectors as ablation values when running circuit localization with EAP. See App. D.2 for details.
@@ 3.3 Results
We present CMD and AUROC scores (Table 2) for each method and task where it is tractable to run; see Table 14 in App. D for CPR scores. On both metrics, EAP-IG-inputs with CF ablations generally performs best. EAP-IG-activations and UGS are competitive with EAP-IG-inputs w.r.t. CMD, but have higher runtime and memory usage respectively. For CPR, they are less competitive; this is unsurprising for UGS, as it directly optimizes for maintaining model behavior, rather than for finding solely positively impactful components.
@@ 4.2 Causal Variable Localization Baselines
We consider five baselines for constructing the featurizer $\mathcal{F}$ and selecting the features $\Pi_{X}$. See App. F for further details.
@@ Results.
We target the residual stream of the last token of the input and the correct choice letter token at each layer. We generally see strong evidence (Figures 3(c) and 3(e); App. F.5.1) of the causal model $\mathcal{H}_{\mathbf{MCQA}}$ being a faithful abstraction, with DAS successfully disentangling the ordering ID variable $X_{\text{Order}}$ from the output token $O_{\text{Answer}}$ in many layers. Even the full vector baseline successfully localizes the variable in some layers, though it performs poorly on average because both variables are aligned with the same features.
Table 3: Baseline results for the causal variable localization track. In each table, the first row is the task, the second row is the model, and the third row is the causal variable. For Arithmetic, MCQA, and ARC (Easy), we report interchange intervention accuracy, i.e., the proportion of aligned interventions on the causal model and deep learning model that result in the same output token(s); higher is better. For each method of aligning a causal variable to LM features, we report the mean across counterfactual datasets and layers in the low-level model. In parenthesis and bold, we report the best alignment across all layers. For IOI, we report the mean-squared error between the causal model logit and the deep learning model logit; lower is better. See App. F.5 for more detailed results by task.
@@
For each baseline, we target the last token and the last token of the second operand. We see poor performance across the board (Figure 3(d); App. F.5.2). The Gemma-2 results are at chance and the Llama-3.1 results have only faint signs of success. The difference in results is likely due to only Llama-3.1 tokenizing multi-digit numbers as single tokens. There may be no “carry-the-one” variable present in these models. Alternatively, the variable might be represented in a non-linear features space, e.g., an onion representation (Csordás et al., 2024). Future submissions that beat baselines will require genuine progress.
@@
For each baseline, we target the last token and the last token of the city entity in the prompt. We generally see evidence that the attributes can be localized and disentangled (Figure 3(g); App. F.5.3), though the Llama-3 results are weaker. Because the dataset requires balancing causing an attribute to change with not changing the other attributes, the full vector baseline entirely fails completely.
@@
First, we replicate these experiments on our curated datasets and fit our own linear model that we use to define a causal model $\mathcal{H}_{\mathbf{IOI}}$ that predicts the logit difference between the indirect object and the subject. The high-level causal model takes in a text input $T$ and computes the token and positional information $S_{\text{Tok}}$ and $S_{\text{Pos}}$ encoding the subject token identity and position, respectively. Then, the output variable $O$: (i) checks if the token and position variables $S_{\text{Tok}}$ and $S_{\text{Pos}}$ match the input $T$ and inverts the signal if a mismatch is detected, and (ii) computes the logit difference between the indirect object and subject as a linear function of these binary variables. See Appendix F.5.4 for details of the causal model.
@@
Broadly, we find evidence that the position and token variable can be disentangled (Figure 3(f); App. F.5.4). For the full vector baseline, we conduct a brute force search and find an alignment of position to heads 7.3, 7.9 ($\Pi_{S_{\text{Pos}}}$), and 8.6 and token to head 8.10 ($\Pi_{S_{\text{Tok}}}$), to be better than other alignments to entire heads. While the variables can be disentangled at the level of heads, even better results are achieved when each variable is aligned to features of heads.
@@ 4.7 General Discussion
DAS is the only method that learns features with supervision from the high-level causal model, so it is not surprising this method performs best. This establishes a skyline to compare unsupervised featurization methods against.
DBM on standard hidden dimensions is a strong baseline.
The success of DBM shows that the dimensions of untransformed hidden vectors can be useful units of analysis; however, DAS outperforming DBM shows that non-basis-aligned directions in activation space are generally better units of analysis.
DBM on PCA or SAE features does not outperform DBM on standard dimensions. SAE and PCA features generally fail to improve upon neurons, i.e., standard dimensions of hidden vectors, as a unit of analysis. Across the board, DBM on SAE features is the worst performing method other than the full vector baselines. A notable exception is the MCQA task where PCA performs very poorly, which we suspect is due to the low-data regime of $\approx$100 examples harming the quality of the principal components we compute. This is in line with the results from the steering benchmark AxBench (Wu et al., 2025), in which SAEs struggle against simple steering baselines.
@@
There are several clear opportunities for future submissions to the causal variable track to improve upon our baselines. First, our baselines provide weak evidence at best for the carry-the-one variable of the two digit addition task and we have hope that more complex feature spaces might be more successful. Second, while we only run baselines for IOI on the GPT-2 attention heads identified by (Wang et al., 2023), we encourage future submissions to locate and featurize attention heads in the other three models. Third, we did not conduct exhaustive hyperparameter searches for any of the baseline methods, so all tasks likely have room for improving the best alignment and average alignment across layers.